OpenStudy (alex_mattucci):

What does it mean to write the expression x^2 + 2x - 24 as a product of two linear expressions?

Nnesha (nnesha):

factor

OpenStudy (alex_mattucci):

So just factor it? That is it?

Nnesha (nnesha):

yep

OpenStudy (alex_mattucci):

I thought all that fancy text was supposed to mean something? All well, that is pretty easy!!!

OpenStudy (alex_mattucci):

Thanks for the help!!!

Nnesha (nnesha):

factor of that equation would be like (x-b)(x+b) which is same as linear equation y=mx+b

Nnesha (nnesha):

np o^_^o

OpenStudy (alex_mattucci):

Does all this fancy text also just mean to factor? Rewrite x^2 - 8x + 16 as a square of a linear expression.

Nnesha (nnesha):

lol i feeeel ya ;-^

Nnesha (nnesha):

yes right there is a word `square` that means u answer would be (x-h)^2 you will get same parentheses here is an example x^2+4x+4 we need two number when we multiply them we should get product of AC =4 times 1 =4 and when we add/subtract them we should get middle which is also 4 in this question so 2 times 2 =4 2+2=4 factors are (x+2)(x+2) which can be written as (x+2)^2

OpenStudy (alex_mattucci):

So would x^2 + 2x - 24 be (x-4)(x+6)?

Nnesha (nnesha):

6 times -4 = -24 yes 6-4 =2 so yes

OpenStudy (alex_mattucci):

And so x^2 - 8x + 16 should be (x-4)^2, correct?

Nnesha (nnesha):

how ? explain plz :=)